# Comparative Analysis of $k$-essence and Quintessence Scalar Field Models: A Data Analysis Approach

### v1

@@ IV Data Analysis
where $k$ is the number of independent parameters present in the model and $N$ represent the total number of data points and $\mathcal{L}_{\rm max}$ represents the maximum likelihood. Performing parameter inference involve careful selection of priors for the model parameters. We use flat prior for all the model parameters and dynamical variables summarized in Tab. [3].
We utilized a well tested and openly accessible Python implementation of the affine-invariant ensemble sampler for Markov Chain Monte Carlo (MCMC), as proposed by Goodman and Weare, known as emcee for the Bayesian analysis. More comprehensive information about emcee and its implementation can be found in ref. [64]. The best-fit values of the model parameters are summarized in Tab. [3]. Our analysis shows that both the quintessence and the $k$-essence models predict $H_{0}\sim 67$ close to the value obtained by Planck2018 collaboration assuming the $\Lambda$CDM model. However, the value of $\Omega_{m_{0}}$ obtained for the $k$-essence model is slightly less than the standard model prediction for both datasets. The evolution of density parameters corresponding to matter and field for Quintessence and $k$-essence models are shown in Figs. [10a, 10b]. The Fig.[9], shows the comparison between the apparent magnitude predicted by the Quintessence, the $k$-essence and the $\Lambda$CDM models. It is evident from the figure that all these models predict almost similar $m(z)$ for the best-fit model parameters and consequently showing almost similar fit for the joint analysis. The model selection based on AIC and BIC shows that $k$-essence model gives slightly better fit to the data as compared to the Quintessence model while the evidence of these models against the $\Lambda$CDM model is weak.
Figure 9: The residue of $m_{\Lambda\rm CDM}-m_{\rm k-essence}$ is plotted in blue, while $m_{\Lambda\rm CDM}-m_{\rm quintessence}$ is shown in dotted pink. These are compared with the data points $\Delta m=m_{\rm Data}-m_{\Lambda\rm CDM}$. The red dashed line represents $\Delta m=m_{\Lambda\rm CDM}-m_{\Lambda\rm CDM}=0$ (the base line). The comparison shows that both the quintessence and $k$-essence models exhibit behavior similar to the $\Lambda$CDM model. All the lines are very near to each other and so it is difficult to differentiate the colors.
@@
Figure 10: Evolution of the (a) quintessence and (b) $k$-essence field corresponding to the best fit values of model parameters obtained from OHD+Pantheon dataset.
@@
V Conclusion
In this study, we analyzed the quintessence and scalar field models with exponential and and inverse square law potentials respectively and compared the theoretical predictions with the $\Lambda$CDM model using the observational probes consisting of pantheon+ compilation of Type Ia supernovae (SNIa) and Observational Hubble data (OHD). We employed Markov Chain Monte Carlo (MCMC) method for the parameter inference. Our analysis demonstrated that quintessence and $k$-essence models are successful in explaining the late time acceleration of the universe. Very few work along this lines, using $k$-essence have ever been done before. Based on the best values obtained from the joint (OHD+Pantheon) analysis, we observe that the evolution of $\Omega_{m}$ in the presence of the $k$-essence field takes slightly lower value than in the quintessence-dominated scalar field, as shown in Figs. [10a, 10b].
We see that in the quintessence-dominated field, during the past epoch, the matter density dominates over the field density, and the corresponding effective equation of state becomes zero, signifying the matter-dominated epoch. However, for the $k$-essence-dominated field, although the matter density dominates, both the field and fluid densities in the past epoch are non-zero and the corresponding effective EoS vanishes, demonstrating that the field may itself behave like some form of matter which produces non-accelerating expansion. Due to this behavior, it can be seen from the evolution plot that in the current epoch, field density is nearly $80\%$, while the matter density is around $20\%$. In all of our work we have used the matter part to be composed of a pressure-less fluid. Apart from the distinct density evolution for both fields, the evolution of the effective equation of state (EoS) produces similar values, yet distinct from the $\Lambda$CDM model at the current epoch. This is our central result. However, for the quintessence field, the effective equation of state appears to stabilize at $-0.50$ in the far future, while for the $k$-essence field, the effective equation of state converges to $-0.52$. The value of $H_{0}$ obtained by the quintessence and $k$-essence models are in close agreement with the value reported by Planck2018 mission assuming the $\Lambda$CDM model. The crucial difference between these models can only be observed with high redshift data such as CMB. According to Akaike and Bayesian information criteria for model selection, the evidence of $\Lambda$CDM model is slightly higher than the Quintessence and $k$-essence scenario.
We have also pointed out that the simplest theory of $k$-essence, only kinetically driven, which can combine both dark matter and dark energy in a combined way cannot be constrained via the observational tools at our hand. The number of parameters in this case make it very difficult for data anlysis. In this paper we primarily focused on models where the scalar field sector was not directly coupled to the fluid sector via any non-minimal coupling. It will be interesting to see how non-minimal coupling in the dark sector modifies the results.
@@ Appendix A A note on pure kinetic $k$-essence
which takes the same form as $\Lambda$CDM. From this, we see that it is not possible to constrain all the model parameters of the kinetic k-essence. One can only obtain constraints on $\Omega_{F_{0}}$, and from this information, one can at best estimate the range of some parameters by assuming certain values for other parameters. Here, we will get the same value for $\Omega_{F_{0}}$ as $\Omega_{\Lambda 0}$ as determined in Tab. 3.

### v2

@@ III Data Analysis
where $k$ is the number of independent parameters present in the model and $N$ represent the total number of data points and $\mathcal{L}_{\rm max}$ represents the maximum likelihood. Performing parameter inference involve careful selection of priors for the model parameters.
Flat $\Lambda$CDM model
Dataset $H_{0}$ $\Omega_{m}$ $r_{d}$ AIC BIC
Priors [30,100] [0,0.6] [100, 300]
CC+BAO $69.8\pm 1.9$ $0.297^{+0.017}_{-0.019}$ $144.1\pm 3.6$ 29.68 34.74
CC+BAO+DESBAO $69.5\pm 1.8$ $0.302^{+0.011}_{-0.013}$ $145.1^{+3.4}_{-3.8}$ 45.11 50.66
CC+BAO+DESBAO+DES $69.93\pm 0.24$ $0.3185\pm 0.0098$ $142.63\pm 0.85$ 1691.39 1708.00
CC+BAO+DESBAO+PP+DES $72.56\pm 0.14$ $0.2691\pm 0.0068$ $142.29\pm 0.85$ 3749.39 3767.94
CC+BAO+DESBAO+PP $73.32\pm 0.17$ $0.315\pm 0.010$ $136.38\pm 0.98$ 1811.50 1827.90
CC+BAO+DES $69.78\pm 0.28$ $0.325\pm 0.012$ $141.7\pm 1.1$ 1675.42 1692.02
CC+BAO+DES+PP $72.73\pm 0.15$ $0.2582\pm 0.0077$ $141.8\pm 1.1$ 3726.06 3744.60
CC+BAO+PP $73.26\pm 0.18$ $0.320\pm 0.012$ $135.5\pm 1.2$ 1795.95 1812.34
CC+DESBAO $69.5\pm 1.8$ $0.304\pm 0.016$ $145.7\pm 3.6$ 31.86 36.85
CC+DESBAO+DES $69.79\pm 0.26$ $0.325\pm 0.012$ $142.7\pm 1.1$ 1676.92 1693.52
CC+DESBAO+DES+PP $72.66\pm 0.15$ $0.2630\pm 0.0074$ $143.6\pm 1.1$ 3732.29 3750.83
CC+DESBAO+PP $73.25\pm 0.18$ $0.320^{+0.011}_{-0.012}$ $136.4\pm 1.3$ 1797.49 1813.88
Table 3: Summary of best-fit values for the parameters at the 68% confidence level for $\Lambda$CDM.
Model I: Inverse Square Potential
Dataset $H_{0}$ $\Omega_{m}$ $r_{d}$ $x_{0}$ $\lambda$ AIC BIC
Priors [30,100] [0,0.8] [100, 300] [0.6, 2.4] [-1.5, 2.0]
CC+BAO $66.9\pm 2.4$ $0.196^{+0.11}_{-0.057}$ $144.7\pm 3.6$ $1.097^{+0.054}_{-0.060}$ $0.34^{+0.19}_{-0.16}$ 30.92 39.36
CC+BAO+DESBAO $67.5\pm 2.1$ $0.185^{+0.12}_{-0.067}$ $145.2\pm 3.7$ $1.125\pm 0.044$ $0.26^{+0.15}_{-0.12}$ 46.28 55.53
CC+BAO+DESBAO+DES $68.99\pm 0.36$ $0.180^{+0.12}_{-0.066}$ $142.76\pm 0.86$ $1.136\pm 0.021$ $0.234^{+0.071}_{-0.062}$ 1685.91 1713.59
CC+BAO+DESBAO+PP+DES $72.84\pm 0.21$ $0.191^{+0.084}_{-0.046}$ $142.07\pm 0.83$ $1.243\pm 0.013$ $-0.051^{+0.054}_{-0.061}$ 3746.29 3777.20
CC+BAO+DESBAO+PP $72.64\pm 0.24$ $0.157\pm 0.090$ $135.72\pm 0.98$ $1.139^{+0.023}_{-0.019}$ $0.251^{+0.066}_{-0.058}$ 1802.39 1829.73
CC+BAO+DES $68.95\pm 0.36$ $0.195^{+0.11}_{-0.052}$ $142.3\pm 1.1$ $1.128\pm 0.021$ $0.255^{+0.072}_{-0.063}$ 1670.43 1698.09
CC+BAO+DES+PP $72.84\pm 0.21$ $0.176^{+0.091}_{-0.049}$ $141.4\pm 1.1$ $1.242\pm 0.014$ $-0.018\pm 0.06$ 3727.72 3758.63
CC+BAO+PP $72.61\pm 0.25$ $0.175^{+0.12}_{-0.059}$ $135.3\pm 1.2$ $1.130\pm 0.023$ $0.275^{+0.071}_{-0.058}$ 1786.94 1814.25
CC+DESBAO $68.8^{+2.4}_{-2.7}$ $0.175^{+0.12}_{-0.073}$ $145.6^{+3.4}_{-3.9}$ $1.173^{+0.052}_{-0.066}$ $0.13^{+0.21}_{-0.15}$ 34.36 42.68
CC+DESBAO+DES $69.09\pm 0.36$ $0.149^{+0.065}_{-0.13}$ $143.2\pm 1.2$ $1.148^{+0.023}_{-0.020}$ $0.221^{+0.076}_{-0.058}$ 1673.30 1700.96
CC+DESBAO+DES+PP $72.86\pm 0.21$ $0.119^{+0.056}_{-0.094}$ $143.1\pm 1.1$ $1.254^{+0.014}_{-0.012}$ $-0.016^{+0.069}_{-0.052}$ 3727.20 3758.10
CC+DESBAO+PP $72.70\pm 0.24$ $0.127^{+0.051}_{-0.12}$ $136.4\pm 1.2$ $1.150^{+0.024}_{-0.020}$ $0.241^{+0.066}_{-0.054}$ 1789.57 1816.88
Table 4: Summary of best-fit values for the parameters at the 68% confidence level for Model I.
Model II: Exponential potential with fixed parameter: $\mathcal{M}_{0}=8$.
Dataset $H_{0}$ $\Omega_{m}$ $r_{d}$ $x_{0}$ $\beta$ AIC BIC
Priors $[30,100]$ [0,0.6] [100, 300] [0.7, 2.2] [-2.5, 2.5]
CC+BAO $66.3\pm 2.5$ $0.187^{+0.12}_{-0.062}$ $144.8\pm 3.7$ $1.071^{+0.032}_{-0.038}$ $-1.25^{+0.52}_{-0.61}$ 31.08 39.52
CC+BAO+DESBAO $67.0\pm 2.1$ $0.166\pm 0.092$ $145.3\pm 3.6$ $1.061\pm 0.029$ $-0.97^{+0.41}_{-0.47}$ 46.21 55.46
CC+BAO+DESBAO+DES $68.93\pm 0.37$ $0.174^{+0.12}_{-0.075}$ $142.79\pm 0.87$ $1.047^{+0.019}_{-0.022}$ $-0.73\pm 0.21$ 1685.90 1713.58
CC+BAO+DESBAO+PP+DES $72.84\pm 0.21$ $0.184^{+0.091}_{-0.051}$ $142.07\pm 0.87$ $1.009^{+0.011}_{-0.017}$ $0.14^{+0.19}_{-0.17}$ 3747.10 3778.01
CC+BAO+DESBAO+PP $72.60\pm 0.25$ $0.151^{+0.079}_{-0.11}$ $135.73\pm 0.97$ $1.052^{+0.020}_{-0.018}$ $-0.78\pm 0.19$ 1802.61 1829.95
CC+BAO+DES $68.92\pm 0.38$ $0.191^{+0.11}_{-0.055}$ $142.3\pm 1.1$ $1.045^{+0.015}_{-0.021}$ $-0.77\pm 0.22$ 1670.49 1698.15
CC+BAO+DES+PP $72.84\pm 0.21$ $0.167^{+0.098}_{-0.055}$ $141.5\pm 1.1$ $1.013^{+0.012}_{-0.017}$ $0.04\pm 0.18$ 3728.46 3759.36
CC+BAO+PP $72.57\pm 0.25$ $0.163^{+0.11}_{-0.088}$ $135.3\pm 1.2$ $1.052^{+0.018}_{-0.020}$ $-0.87^{+0.19}_{-0.22}$ 1787.15 1814.46
CC+DESBAO $67.8\pm 2.6$ $0.153\pm 0.093$ $145.7^{+3.4}_{-3.8}$ $1.051\pm 0.034$ $-0.69^{+0.53}_{-0.63}$ 34.19 42.50
CC+DESBAO+DES $69.02\pm 0.37$ $0.141^{+0.064}_{-0.12}$ $143.3\pm 1.1$ $1.051^{+0.022}_{-0.016}$ $-0.70^{+0.19}_{-0.23}$ 1673.09 1700.75
CC+DESBAO+DES+PP $72.87\pm 0.21$ $0.113^{+0.047}_{-0.098}$ $143.1\pm 1.1$ $1.020^{+0.018}_{-0.010}$ $0.04^{+0.16}_{-0.21}$ 3727.83 3758.73
CC+DESBAO+PP $72.65\pm 0.25$ $0.116^{+0.044}_{-0.11}$ $136.4\pm 1.2$ $1.056^{+0.022}_{-0.012}$ $-0.77^{+0.17}_{-0.21}$ 1789.38 1816.69
Table 5: Summary of best-fit values for the parameters at the 68% confidence level for Model II.
@@
Figure 6: Evolution of the $k$-essence field for (a) Inverse square potential and (b) exponential potential corresponding to the best fit values of model parameters obtained from the data analysis.
We utilized a well tested and openly accessible Python implementation of the affine-invariant ensemble sampler for Markov Chain Monte Carlo (MCMC), as proposed by Goodman and Weare, known as emcee for the Bayesian analysis. More comprehensive information about emcee and its implementation can be found in ref. [72].
Marginalized 1D and 2D posterior distribution of the dynamical variables and model parameters for the $\Lambda$CDM model, Model I and Model II for various data sets are shown in Fig. [7]. The mean and standard deviation of the distribution are considered as the best-fit value and error corresponding to each model parameters, which are summarized in Tabs. [3, 4, 5] respectively. In the table, we provided the dynamical variables and model parameters corresponding to Model I and Model II and specified the corresponding prior ranges. These prior ranges were selected based on the analysis conducted in the previous dynamical stability section. Note that the parameter ranges specified in Eqs. (18) and (33) are extended in Tabs. [4, 5].
For Model I, which refers to the $k$-essence field with inverse square potential, there are mainly two primary variables, $(x,y)$, whose initial conditions $(x_{0},y_{0})$ must be varied. However, in the table, we have listed $x_{0}$ and $\Omega_{m0}$ instead of $y_{0}$. The reason for this is clarified in Eq. (21). On the other hand, for Model II, we have followed the same approach, constraining all the necessary parameters except $\mathcal{M}_{0}$. The parameter $\mathcal{M}_{0}$ is fixed to a certain value $\mathcal{M}_{0}=8.0$, as this parameter must be a positive number. We emphasize that although this parameter is fixed, the field variables $(x_{0},y_{0})$ are still varying, as can be seen from the Hubble equation Eq. (27), which reads as:
$y_{0}=\frac{\log\left(-\frac{12(\Omega_{m0}-1)}{\mathcal{M}_{0}\left(3x^{4}-2x^{2}\right)}\right)}{\beta}.$ (48)
Our analysis shows that both models predict the value of $H_{0}$ in close agreement with each other, but with a significant deviation from the $\Lambda$CDM model. The present-time matter energy density $\Omega_{m0}$ is predicted to be lower for all datasets in both models compared to $\Lambda$CDM.
The evolution of cosmological parameters in Model I for the best-fit values provided in Tab. [4], such as the fractional energy densities and the effective equation of state, is shown in Fig. [6a]. In this plot, the evolution is depicted against the number of e-folds $N$. The results demonstrate that the model exhibits an extended matter-dominated phase (i.e., $\Omega_{m}>\Omega_{\phi}$), for all datasets, before transitioning to an accelerating epoch. During this matter-dominated phase, the effective equation of state is nearly zero, indicating attractive scaling behavior. As the system enters the late-time phase, the model shows phantom behavior for CC+BAO+DESBAO+PP+DES data, $\omega_{\rm eff}\sim-1.05$ and accelerating behavior for rest of the datasets.
Figure 7: 1D and 2D marginalized posterior distribution of the model parameters of (a) the flat $\Lambda$CDM and the $k$-essence field with (b) inverse square type and (c) exponential type potential with fixed $\mathcal{M}_{0}=8$ for different datasets.
The cosmological parameters of Model II, corresponding to the best-fit values, are plotted against $\tilde{N}$ in Fig. [6b], as the autonomous system of equations is regulated at the coordinate $z=-1$. In this case, a matter-dominated phase persists for an extended period during the past epoch in terms of e-folds $\tilde{N}$. As the system enters the current epoch, the field density increases to 70%, leading to acceleration. In the future epoch, an interesting scenario arises where, the model exhibits an accelerating solution, but the effective EoS begins to increase towards positive values, indicating a phase transition from an accelerating to a non-accelerating scenario. On the other hand, for the CC+BAO+DESBAO+PP+DES, the model exhibits stable phantom behavior $\omega_{\rm eff}\to-1$.
Additionally, we evaluated the information criteria, specifically AIC and BIC, for the models. For this comparison, $\Lambda$CDM was considered the base model. The model having lower value of AIC or BIC is preferred over the other. From Tab. [3], [4] and [5], it is evident from the AIC criterion that Model 1 and Model 2 are preferred over the $\Lambda$CDM model for most data combinations. However, according to the BIC, the $\Lambda$CDM model is preferred over Model I and Model II for all data combinations, as BIC more effectively penalizes models with extra number of parameters.
@@
IV Conclusion
[…]
Many studies have been conducted with the inverse square potential, but few have analyzed the exponential type potential with this particular form of the Lagrangian. One key challenge in the dynamics of the exponential potential is the need for more than two dynamical variables to close the autonomous system of equations, resulting in a 3D phase space. Due to the different structure of dynamical variables compared to the inverse square potential, the exponential model does not yield critical points that signify the current accelerating or phantom phase. The critical points only describe either far past or future epochs. Thus, model parameters such as $\beta$ and $\mathcal{M}_{0}$ could not be constrained with the arguments presented for Model I. Therefore, we relied on numerical evolution and skipped the critical point analysis, varying the initial conditions and model parameters within a specific range to see whether the model produces different phases of the universe. We found that for certain ranges, the model produces a matter-dominated phase, and as the system enters the current epoch, the system’s equation of state becomes lower than the background fluid’s EoS, exhibiting tracking behavior. Thus, the exponential model can produce a stable accelerating (phantom) phase followed by an attractive matter phase for certain ranges, alleviating the fine-tuning issue.
However, the picture becomes clearer after employing the Markov Chain Monte Carlo (MCMC) method for parameter inference. The analysis demonstrated that the inverse square type model can produce a late-time accelerating scenario for all datasets. Notably, for certain datasets, the model exhibits a phantom behavior with $\omega_{\rm eff}<-1$ in the future epoch, indicating a significant deviation from the concordance $\Lambda$CDM model.
On the other hand, Model II, which corresponds to the exponential type potential, displays a similar scenario to Model I, where the matter phase dominates for a certain period, and the matter energy density exceeds the field energy density. However, an intriguing aspect of this model emerges: in the late-time phase, for the combined datasets, the model demonstrates a cosmological constant-type behavior, persisting over an extended period. Additionally, for the other datasets, the model undergoes a phase transition in the future epoch, shifting from an accelerating phase to a non-accelerating phase. This significant difference adds complexity to the model’s dynamics and suggests that further studies, particularly with Planck datasets, are necessary to place stringent constraints on the model parameters.
We further assessed the viability of the models using statistical criteria, specifically the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), and compared the results with those of the $\Lambda$CDM model. For most of the combined datasets, both models outperformed $\Lambda$CDM according to AIC criterion. However, $\Lambda$CDM remained more favorable for all the data combinations when we considered BIC criterion. This is because BIC criterion penalizes models with additional parameters. This suggests that while the current models provide good fit to the all the data combinations, the $\Lambda$CDM model is preffered as compared to k-essence models.
Acknowledgements.
One of the authors Sarath Nelleri is thankful to IIT Kanpur for providing with the institute postdoctoral fellowship. We would like to acknowledge the use of PARAM Sanganak, a high-performance computing facility at the Indian Institute of Technology Kanpur (IITK), for carrying out the data analysis work.
@@ Appendix A A note on pure kinetic $k$-essence
which takes the same form as $\Lambda$CDM. From this, we see that it is not possible to constrain all the model parameters of the kinetic k-essence. One can only obtain constraints on $\Omega_{F_{0}}$, and from this information, one can at best estimate the range of some parameters by assuming certain values for other parameters. Here, we will get the same value for $\Omega_{F_{0}}$ as $\Omega_{\Lambda 0}$ as determined in Tab. [3].